1.two identical balls are shot upwards one after the other at an interval 2 sec along the same vertical line with same initial velocity of 40 m/s. What is the height at which balls collide?


a)50m b)75m c)100m d)125m


 


2. A football is kicked vertically upward from the ground and a person gazing out of window sees it moving upwards past him at 5 m/s.Window is 15m above the ground.Ignore air resistance And take g=10.


How high does the ball go above ground? and how much time does it take to go from ground to highest point?


 


3. Can you tell me name of any publisher or author for solutions of HC Verma and DC Pandey?

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chilukuri

let after t sec of the second shot ball will colide

for first ball

h1=40*(t+2) -1/2 g (t+2)2

for second ball

h2=40* t -1/2 g t2

for collison

h1=h2

40*(t+2) -1/2 g (t+2)2  =40* t -1/2 g t2

40=g(t+1)

now put value of t in any one of the above equation u will get h=75 m


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

 

Palak
5 years ago
Two balls of massess are thrown upward along the same vertacal line at an interval of 2 seconds , with the same initial velocity of 40m/sec.then this cillide at the height of(Take g = 10 m/second square)